# Zhou, Orfanoudaki, and Zhu

### v2

@@ 7.2 Decision Prescription Evaluation
Evaluated empirical confidence ranking ($\downarrow$) for different methods across three datasets.
@@ 7.3 Ablation Studies
Comparison of three different calibrated radii.
@@
Comparison of four different configurations of the generative models. Point is constant across different specifications of $K$ and $\sigma$, as they do not affect the fitting of the point prediction model.
@@
For the second set of ablation studies, we implement CREDO with four different configurations of the underlying prediction (generative) model: ($i$) Oracle: using the ground-truth data distribution as the generative model. ($ii$) Generative: a three-component Gaussian Mixture model, fitted using the EM algorithm for $1\times 10^{2}$ epochs. (iiiiii) Generative (perturbed): same as Generative except that the fitted mean of the components is perturbed with a small noise, which follows a $U(0.5,1)$ distribution. ($iv$) Point: point prediction model that captures the mean of the marginal distribution of $Y$. The metrics used in this experiment follow those in the risk-estimation evaluation (Section 7.1), with the addition of the true positive rate (TPR).
@@
In the last set of experiments, we evaluate the performance of different optimization procedures solving Problem (14b). We use three general metrics: ($i$) Obj: the solution’s objective value averaged across trials; ($ii$) Vio: the average ratio of solutions that violate the optimization constraint. and (iiiiii) Err: the average distance between the solution and the ground-truth solution. The ground truth solution is computed from the closed-form solution under the linear setting, and approximated using a highly granular brute-force enumeration strategy in nonlinear settings. We include five baselines: ($i$) GD: a generic gradient descent algorithm solving the optimization by penalizing the objective function with the constraint; ($ii$) BF: brute force enumeration, where its discretization resolution is set to be the same as the number of iterations $T$ for other methods; (iiiiii) RS: random search algorithm, which is similar to BF except that the enumeration is done via sampling from a standard Gaussian distribution; ($iv$) RG: random guess, which is equivalent to a one-step RS. These baselines are chosen as they are to the best of our knowledge, the few tractable procedures that can effectively handle Problem (14b). To ensure fair comparison, all the above optimization baselines are tuned with the number of iterations and/or epochs such that they take approximately the same time to run as our proposed method.
@@ 8 Discussion and Conclusions
Our theoretical analysis establishes that CREDO achieves conservative risk estimates, providing expectation-wise validity even under model misspecification. The framework’s consistency properties ensure that these bounds tighten asymptotically as more data becomes available, while the true positive rate analysis demonstrates that generative sampling effectively reduces excessive conservatism. These properties position CREDO as both theoretically rigorous and practically useful, addressing the long-standing challenge of quantifying decision stability without distributional assumptions.
@@ 10 Proofs
In this section, we denote $k$-th prediction set constructed by $\hat{y}^{(k)}$ as $\mathcal{C}^{k}(\cdot;\cdot)$ for notation clarity.
@@
Denote the $k$-th prediction set constructed by $\hat{y}^{(k)}$ as $\mathcal{C}^{k}$. Then under Section 4.1,
@@
Proof 10.2
Proof of Lemma 10.1 When $\hat{\alpha}<1/(n+1)$ or $\hat{\alpha}=1$, by the definition of all three variants of $R(\alpha)$, the statement holds trivially. So we only need to consider the case when $1>\hat{\alpha}\geq 1/(n+1)$. For any $k=1,\ldots,K$, we begin by expanding the left-hand side:
@@
It can be seen that the following two conditions hold:
@@
Using a first-order Taylor expansion on the left-hand side of the inequality above, there is
@@
We assume that this approximation is exact, so the $\approx$ sign can be replaced with a == sign. Consequently, by combining the two equations above, we get
@@
The first equality follows from our problem reformulation (Lemma 4.1). The second inequality holds due to the definition of $\tilde{\alpha}^{(k)}(z)$, which guarantees that the $k$-th generated CP region is always contained in $\pi^{-1}(z)$. Since by Lemma 10.1, we have proved that:
@@
We conclude the entire proof for Theorem 5.1.
Remark 10.4 (Discussion on the Taylor approximation error)
We comment on the approximation error of the first-order Taylor approximation in (21). This approximation trick has been adopted in prior works (gauthier2025values), which has been argued that its error is small when the estimator $\hat{\alpha}$ is well concentrated around its mean. Empirically, this condition is usually satisfied in our setting. For example, when CREDO is deployed in a human-algorithm collaboration setting, the candidate decisions provided from the decision maker would be expected to be near optimal and should already enjoy a relatively small ground truth risk. This makes $\hat{\alpha}$ have a relatively small variance and well concentrated around its mean.
@@
Then, plugging them into the previous equation, we get
@@
to achieve a tighter bound (i.e., smaller offset). One can prove that as long as $\delta$ is chosen such that $\delta>1/(n+1)$, all theorems presented in the main text remain valid, and the bound (i.e., offset) becomes tighter as $\delta$ increases. Then, one can adopt $R^{\prime}(\alpha)$ as the conformalized radius and take (23) as the final estimator. This procedure is equivalent to truncating the lower part of $\hat{\alpha}(z)$ at $\delta$, i.e., setting $\max\{\hat{\alpha}(z),\delta\}$ as the risk estimator, and then taking (23) as the final estimator, which is simpler to execute in implementation.
@@
Proof 10.5
Proof
The proof mainly follows the proof of Theorem 1 in barber2023conformal, with the exception that all steps are conditioned on the $\sigma$-field generated by $\tilde{\alpha}^{(k)}(z)$ [^4] We note that the the $\sigma$-field generated by $\tilde{\alpha}^{(k)}(z)$ is the same across all $k$, as they are just different simulation trials based on the same input and generative model, therefore they are equal to the $\sigma$-field of $\hat{\alpha}(z)$.. Following the proof of Theorem 1 in barber2023conformal, we can derive that:
@@
Plugging this equation back in (22), we get
$\mathbb{P}\left\{\pi(Y)=z\right\}\geq 1-\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\left[\tilde{\alpha}^{(k)}(z)+\frac{\sum_{i=1}^{n}\mathbb{E}\left[d_{\rm TV}^{(i)}(\hat{\alpha}(z))\right]}{n+1}\right]=1-\mathbb{E}\left[\tilde{\alpha}(z)\right]-\frac{\sum_{i=1}^{n}d_{\rm TV}^{(i)}(\hat{\alpha}(z))}{n+1}.$
@@
Proof 10.6
Proof By Remark 6.3,
@@
Therefore, for any $x\in\mathcal{X}$, conditioning on $X=x$, there is
@@
[…]
$=\frac{1}{K}\sum_{k=1}^{K}T_{k}^{(1)}+\frac{1}{K}\sum_{k=1}^{K}T_{k}^{(2)}.$
For the first term, since $T_{k}^{(1)}$ are i.i.d. random variables [^5] Since $\hat{y}^{(k)}$ are independently generated from the same generative model given $x$., and $\mathbb{E}[T_{k}^{(1)}\mid X=x]=0$, assuming that its variance is finite, by the central limit theorem, there is $\frac{1}{K}\sum_{k=1}^{K}T_{k}^{(1)}=O_{p}(K^{-1/2})$. For the second term, note that by the assumption
[…]
@@
$A$ $=\{z\in\mathcal{Z}\mid\mathbb{P}\{z\not\in\pi(Y)\}<1\text{ and }\hat{\alpha}(z)<1\},$
$B$ $=\{z\in\mathcal{Z}\mid\mathbb{P}\{z\not\in\pi(Y)\}<1\}.$
Note that $A\subseteq B$. Without loss of generality, we assume that $B$ (therefore $A$) is a finite set, i.e., there is only a finite set of decisions that have ground-truth risk smaller than one [^6] This proof naturally extends to the infinite case by replacing the counting measure “#” with continuous measures, such as Lebesgue measure defined within the decision space $\mathcal{Z}$.. We also define
$w^{(k)}\coloneqq 1-R^{-1}\left(\min_{y\in\partial\pi^{-1}(z)}\|\hat{y}^{(k)}-y\|_{2}\right).$
Since $B$ is a constant irrelevant to $K$, we begin by expanding the following expression:
$\mathbb{E}\left[\#A\right]$ $=\mathbb{E}\left[\sum_{z\in B}\mathbbm{1}\left\{\hat{\alpha}(z)<1\right\}\right]$
$=\mathbb{E}\left[\sum_{z\in B}\left(1-\prod_{k=1}^{K}\left(\mathbbm{1}\left\{\hat{y}^{(k)}_{n+1}\notin\pi^{-1}(z)\right\}\cdot\mathbbm{1}\left\{w^{(k)}>0\right\}\right)\right)\right]$
$=\mathbb{E}\left[\#B\right]-\mathbb{E}\left[\sum_{z\in B}\prod_{k=1}^{K}A_{k}\right].$
Here, the second equality results from the observation that when: ($i$) at least one of the model predictions falls within $\pi^{-1}(z)$, and ($ii$) its calibrated weight is not zero, then the estimated risk after averaging would be less than one. In the last equality, we denote
$A_{k}=\mathbbm{1}\left\{\hat{y}^{(k)}\notin\pi^{-1}(z)\right\}\cdot\mathbbm{1}\left\{w^{(k)}>0\right\}.$
Note that $A_{k}\in\{0,1\}$ and has a nonzero probability taking one [^7] Under some regularity condition on the distribution of $Y$, e.g., is supported on the whole space $\mathcal{Y}$. Without this condition, the statement also holds by changing “monotonically increasing” to “non-decreasing”.. Since at the right hand side, the term $\prod_{k=1}^{K}A_{k}$ monotonically decreases with $K$ almost surely, then $\mathbb{E}\left[\#A\right]$ monotonically increase with $K$. Since $\text{TPR}=\mathbb{E}\left[\#A\right]/\mathbb{E}\left[\#B\right]$, we know that TPR monotonically increases with $K$. This concludes the proof.
@@
Proof 10.8
Proof We begin by noticing that the set containment condition $\mathcal{C}(x;\alpha)\subseteq\pi^{-1}(z)$ is equivalent to requiring that the calibrated conformal ball does not intersect the complement of the inverse feasible region:
@@
so the first constraint in (25) can be replaced as
$\min_{\alpha}\max_{y}\alpha\quad\text{s.t.}\quad\alpha\geq R^{-1}\left(\|y-\hat{y}\|_{2}\right)\quad\text{and}\quad y\in\bigl(\pi^{-1}(z)\bigr)^{c},$
which makes $\alpha$ a slack variable that can be dropped by moving the constraint expression to the objective:
@@
Finally, by definition of $\pi^{-1}$, the constraint can be expanded as
@@
10.6 Proof of Proposition 6.4
Proof 10.9
Proof We begin the derivation by expanding the indicator term:
$\mathbbm{1}\left\{y\in\pi^{-1}(z)\right\}$ $=\mathbbm{1}\left\{f(y,z)\leq\min_{z^{\prime}\in\mathcal{Z}}f(y,z^{\prime};\theta)+\epsilon\right\}$
$=\prod_{z^{\prime}\in\mathcal{Z}}\mathbbm{1}\left\{f(y,z)\leq f(y,z^{\prime})+\epsilon\right\}$
$=\prod_{z^{\prime}\in\mathcal{Z}}\mathbbm{1}\left\{\langle y,z-z^{\prime}\rangle\leq\epsilon\right\}$
$=\mathbbm{1}\left\{\max_{z^{\prime}\in\mathcal{Z}}\langle y,z-z^{\prime}\rangle\leq\epsilon\right\}$
@@
$\max_{z^{\prime}\in\mathcal{Z}}\langle y,z-z^{\prime}\rangle=\max_{v\in\mathcal{V}}\langle y,z-v\rangle.$
Plugging this into the tight hand side of the derivation above, we get
@@
Therefore, we conclude that
$\mathbbm{1}\left\{\hat{y}^{(k)}\in\pi^{-1}(z)\right\}=\prod_{v\in\mathcal{V}}\mathbbm{1}\left\{\langle\hat{y}^{(k)},z-v\rangle\leq 0\right\}.$
Combining the above statement with Proposition 6.1, (19) can be expanded as
$\min_{y}\|y-\hat{y}\|_{2}\quad\text{s.t.}\quad\max_{v\in\mathcal{V}}\langle\hat{y},z-v\rangle\leq 0$
[…]
11 Experiment Details
This section is organized as follows: Section 11.1 shows the computation resources used in our experiments. Section 11.2 and Section 11.3 presents the detailed configurations for the two synthetic settings and the real-world case study, respectively. Section 11.4 describes the detailed configurations of CREDO, the baselines, and the procedures for the two experiments (Conservativeness versus Accuracy Tradeoff and Decision Quality Evaluation).
11.1 Computation Resources
Table 1: Computation resources specfications
Operating System CPU RAM GPU Windows 11 13th-generation Intel Core i7, 16 cores 16GB Not used

### v3

@@ 7.2 Decision Prescription Evaluation
Evaluated empirical confidence ranking ($\downarrow$) for different methods across different optimization settings.
@@ 7.3 Ablation Studies
Comparison of CREDO performance for three different calibrated radii.
@@
Comparison of CREDO performance for four different generative model configurations. Point shows constant trend across different specifications of $K$ and $\sigma$, as they do not affect the fitting of the point model.
@@
For the second set of ablation studies, we implement CREDO with four different configurations of the underlying prediction (generative) model: ($i$) Oracle: using the ground-truth data distribution as the generative model. ($ii$) Generative: a three-component Gaussian Mixture model, fitted using the EM algorithm for $1\times 10^{2}$ epochs. (iiiiii) Generative (perturbed): same as Generative except that the fitted mean of the components is perturbed with a small noise, which follows a $U(0.5,1)$ distribution. ($iv$) Point: point prediction model that captures the mean of the marginal distribution of $Y$. The metrics used in this experiment follow those in the risk-estimation evaluation (Section 7.1), with additional inclusion of the true positive rate (TPR) in (12).
@@
In the third set of experiments, we evaluate the performance of different optimization procedures solving Problem (13b). We use three general metrics: ($i$) Obj: the solution’s objective value averaged across trials; ($ii$) Vio: the average ratio of solutions that violate the optimization constraint. and (iiiiii) Err: the average distance between the solution and the ground-truth solution. The ground truth solution is computed from the closed-form solution under the linear setting, and approximated using a highly granular brute-force enumeration strategy in nonlinear settings. We include five baselines: ($i$) GD: a generic gradient descent algorithm solving the optimization by penalizing the objective function with the constraint; ($ii$) BF: brute force enumeration, where its discretization resolution is set to be the same as the number of iterations $T$ for other methods; (iiiiii) RS: random search algorithm, which is similar to BF except that the enumeration is done via sampling from a standard Gaussian distribution; ($iv$) RG: random guess, which is equivalent to a one-step RS. These baselines are chosen as they are to the best of our knowledge, the few tractable procedures that can effectively handle Problem (13b). To ensure fair comparison, all the above optimization baselines are tuned with the number of iterations and/or epochs such that they take approximately the same time to run as our proposed method.
@@ 8 Discussion and Conclusions
Our theoretical analysis establishes that CREDO achieves conservative risk estimates, providing expectation-wise validity even under model misspecification. The framework’s consistency properties ensure that these bounds tighten asymptotically as more data becomes available, while the true positive rate analysis demonstrates that generative sampling effectively reduces excessive conservativeness. These properties position CREDO as both theoretically rigorous and practically useful, addressing the long-standing challenge of quantifying decision stability without distributional assumptions.
@@ 9 Proofs
For notation clarity, we will denote the $k$-th prediction set constructed by $\hat{y}^{(k)}$ as $\mathcal{C}^{k}(\cdot;\cdot)$ in this and the following sections. We begin by proving E-value post-hoc validity, which is a key lemma to proving Theorem 5.1.
@@
Denote the prediction set constructed by $\hat{y}^{(k)}$ as $\mathcal{C}^{k}$. Then under Section 4.1,
@@
Proof 9.2
Proof of Lemma 9.1 When $\hat{\alpha}<1/(n+1)$ or $\hat{\alpha}=1$, by the definition of all three variants of $R(\alpha)$, the statement holds trivially. So we only need to consider the case when $1>\hat{\alpha}\geq 1/(n+1)$. Recall that $\hat{y}_{i}$ denotes a single prediction generated from the calibration data $g(x_{i})$. For any $k=1,\ldots,K$, we begin by expanding the left-hand side:
@@
One can prove that the following two conditions hold:
@@
Using a first-order Taylor expansion, the left-hand side of the first inequality is:
@@
Assuming that this approximation is exact (see discussion in Remark 9.4), and by combining the two equations above, there is
@@
The first equality follows from our problem reformulation (Lemma 4.1). The second inequality holds due to the definition of $\tilde{\alpha}^{(k)}(z)$, which guarantees that the $k$-th generated CP region is always contained in $\pi^{-1}(z)$. By Lemma 9.1, there is:
@@
[…]
Remark 9.4 (Discussion on the Taylor approximation)
We comment on the rationality of the first-order Taylor approximation used in (19). To begin with, the approximation trick has been adopted in prior works (gauthier2025values), and argued that its error is small when the estimator $\hat{\alpha}$ is well concentrated around its mean. Such a condition is usually satisfied in our setting. For example, when CREDO is deployed in a human-algorithm collaboration setting, the candidate decisions provided from the decision maker would be expected to be near optimal and should already enjoy a relatively small ground truth risk. This makes $\hat{\alpha}$ have a relatively small variance and well concentrated around its mean, which allows the Taylor approximation to be relatively tight.
@@
Plugging them into the first equation, we get
@@
to achieve a tighter bound (i.e., smaller offset). One can prove that as long as $\delta$ is chosen such that $\delta>1/(n+1)$, all theorems presented in the main text remain valid. So one can safely adopt $R^{\prime}(\alpha)$ as the conformalized radius and take (21) as the final estimator. In the meantime, this procedure can be equivalently viewed as truncating the lower part of $\hat{\alpha}(z)$ at $\delta$, i.e., setting $\max\{\hat{\alpha}(z),\delta\}$ as the risk estimator, and then taking (21) as the final estimator.
@@
Proof 9.5
Proof Following the proof of Theorem 1 in barber2023conformal, with the exception that all steps are conditioned on the $\sigma$-field generated by $\tilde{\alpha}^{(k)}(z)$, there is:
@@
Here we implicitly utilized the fact that the $\sigma$-field generated by $\tilde{\alpha}^{(k)}(z)$ is the same across all $k$, as they are just different simulation trials based on the same input and generative model, therefore they are equal to the $\sigma$-field of $\hat{\alpha}(z)$. Plugging this equation back in (20), we get
$\mathbb{P}\left\{z\in\pi(Y)\right\}\geq 1-\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\left[\tilde{\alpha}^{(k)}(z)+\frac{\sum_{i=1}^{n}\mathbb{E}\left[d_{\rm TV}^{(i)}(\hat{\alpha}(z))\right]}{n+1}\right]=1-\mathbb{E}\left[\tilde{\alpha}(z)\right]-\frac{\sum_{i=1}^{n}d_{\rm TV}^{(i)}(\hat{\alpha}(z))}{n+1}.$
@@
Proof 9.6
Proof By Remark 6.3, when using $R_{\infty}$, the risk estimator is a Monte Carlo estimator:
@@
Therefore, for any $x\in\mathcal{X}$, conditioning on $X=x$ and given $f(x)=\hat{\mathcal{P}}_{Y|X=x}$, there is
@@
$=\frac{1}{K}\sum_{k=1}^{K}\left(\mathbbm{1}\left\{\hat{y}^{(k)}\in\pi^{-1}(z)\right\}-\mathbb{P}\{z\in\pi(Y)\mid Y\sim\mathcal{P}_{Y|X=x}\}\right)$
$=\frac{1}{K}\sum_{k=1}^{K}\Big(\underbrace{\mathbbm{1}\left\{\hat{y}^{(k)}\in\pi^{-1}(z)\right\}-\mathbb{P}\{z\in\pi(Y)\mid Y\sim\hat{\mathcal{P}}_{Y|X=x}\}}_{A_{k}}\Big)+$
$\quad\frac{1}{K}\sum_{k=1}^{K}\Big(\underbrace{\mathbb{P}\{z\in\pi(Y)\mid Y\sim\hat{\mathcal{P}}_{Y|X=x}\}-\mathbb{P}\{z\in\pi(Y)\mid Y\sim\mathcal{P}_{Y|X=x}\}}_{B_{k}}\Big)$
$=\frac{1}{K}\sum_{k=1}^{K}A_{k}+\frac{1}{K}\sum_{k=1}^{K}B_{k}.$
For the first term, since: ($i$) $A_{k}$ are i.i.d. random variables, as $\hat{y}^{(k)}$ are independently generated from the same generative model given $x$; (iiii) $\mathbb{E}[A_{k}\mid X=x]=0$, (iiiiii) $A_{k}$ has finite variance, therefore by the central limit theorem, there is $\frac{1}{K}\sum_{k=1}^{K}A_{k}=O_{p}(K^{-1/2})$. For the second term, note that by the assumption
$T_{k}^{(2)}=\mathbb{P}\{Y\in\pi^{-1}(z)\mid Y\sim\hat{\mathcal{P}}_{Y|X=x}\}-\mathbb{P}\{Y\in\pi^{-1}(z)\mid Y\sim\mathcal{P}_{Y|X=x}\}\leq d_{\rm TV}(\hat{\mathcal{P}}_{Y|X=x},\mathcal{P}_{Y|X=x})\leq\delta,$
@@
$A$ $=\{z\in\mathcal{Z}\mid\alpha(z)<1\text{ and }\hat{\alpha}(z)<1\},$
$B$ $=\{z\in\mathcal{Z}\mid\alpha(z)<1\}.$
For simplicity, we assume that $B$ (therefore $A$) is a finite set, i.e., there is only a finite set of decisions that have ground-truth risk smaller than one (though note that Proposition 5.4 and its proof naturally extend to the infinite case by replacing the counting measure “#” with continuous measures, such as Lebesgue measure defined within the decision space $\mathcal{Z}$). Since $B$ is irrelevant to $K$, we begin by expanding the following expression:
$\mathbb{E}\left[\#A\right]$ $=\mathbb{E}\left[\sum_{z\in\mathcal{Z}}\mathbbm{1}\left\{\hat{\alpha}(z)<1~\text{and}~\alpha(z)<1\right\}\right]$
$=\mathbb{E}\left[\sum_{z\in B}\mathbbm{1}\left\{\hat{\alpha}(z)<1\right\}\right]$
$=\mathbb{E}\left[\sum_{z\in B}\left(1-\prod_{k=1}^{K}\left(\mathbbm{1}\left\{\hat{y}^{(k)}\notin\pi^{-1}(z)\right\}\right)\right)\right]$
[…]
@@
Proof 9.8
Proof The first part of the proof resembles the main text. We begin by noticing that the set containment condition $\mathcal{C}(x;\alpha)\subseteq\pi^{-1}(z)$ is equivalent to requiring that the calibrated conformal ball does not intersect the complement of the inverse feasible region:
@@
so the first constraint in (23) can be replaced, and the optimization can be rewritten as
$\min_{\alpha}\max_{y}\alpha\quad\text{s.t.}\quad\alpha\geq R^{-1}\left(\|y-\hat{y}\|_{2}\right)\quad\text{and}\quad y\in\bigl(\pi^{-1}(z)\bigr)^{c}.$
In this problem. $\alpha$ can be viewed as a slack variable, and can be dropped so that the constraint becomes the objective:
@@
This turns the original minmax problem into a single-level maximization problem. Finally, by definition of $\pi^{-1}$, the constraint can be expanded as
@@
9.6 Proof of Corollary 6.4
Proof 9.9
Proof Using the linear assumption, we begin the derivation by expanding the constraint:
$f(z;y)>\min_{z^{\prime}\in\mathcal{Z}}f(z^{\prime};y)+\epsilon~\Longleftrightarrow~\min_{z^{\prime}\in\mathcal{Z}}\langle y,z^{\prime}-z\rangle+\epsilon<0$
@@
$\min_{z^{\prime}\in\mathcal{Z}}\langle y,z^{\prime}-z\rangle=\min_{v\in\mathcal{V}}\langle y,v-z\rangle.$
Plugging this into (13b) derived in Proposition 6.1, and using similar reformulation arguement as (18), we only need to solve the following optimization problem:
$\min_{y}\|y-\hat{y}\|_{2}\quad\text{s.t.}\quad\exists v\in\mathcal{V}\setminus\{z\},\ \langle y,v-z\rangle+\epsilon<0$
This optimization finds the closest distance from a point $\hat{y}$ to any halfspaces defined by norm vectors zvz-v and offset $\epsilon$. By using the well-known point to halfspace distance formula, the optimal objective value of this optimization problem is:
$\min_{v\in\mathcal{V}\setminus\{z\}}\frac{|\langle\hat{y},z-v\rangle-\epsilon|}{\|z-v\|_{2}}.$
Therefore, the expression for the risk estimate is
[…]
Remark 9.10 (Indicator term)
We note that under the linear case, the indicator term also admits the following closed-form:
$\mathbbm{1}\left\{\hat{y}^{(k)}\in\pi^{-1}(z)\right\}=\prod_{v\in\mathcal{V}}\mathbbm{1}\left\{\langle\hat{y}^{(k)},z-v\rangle\leq 0\right\}.$ (24)
To see why, we begin with the following expansion:
$\mathbbm{1}\left\{y\in\pi^{-1}(z)\right\}=\mathbbm{1}\left\{f(z;y)\leq\min_{z^{\prime}\in\mathcal{Z}}f(z^{\prime};y)+\epsilon\right\}=\mathbbm{1}\left\{\max_{z^{\prime}\in\mathcal{Z}}\langle y,z-z^{\prime}\rangle\leq\epsilon\right\}.$
Using the Krein–Milman theorem from the proof above, we can derive that the right-hand side is equal to:
@@
Finally, setting $y=\hat{y}^{(k)}$ yields (24), which concludes the proof.
10 Additional Experiment Details
This section presents additional details of the numerical experiments that extend what has been described in the main text, and is organized as follows: Section 10.1 shows the resources and dependencies; Section 10.2 presents the detailed configurations for the optimization problems; Section 10.3 describes the detailed configurations of the used baselines; Section 10.4 and Section 10.5 describes the metrics and additional experiment results.
10.1 Implementation Environments